1.What is the vapor pressure of the pure solvent if the vapor pressure of a solution of 15g sucrose (C12H22O11) in 100g of ethanol (C2H5OH) is 58 mmHg?


2.A solution prepared by adding 10g of an unknown non-volatile, non-electrolyte solute to 150 g of water. The Boiling point of the solution is elevated by 0.433 0C above the normal boiling point of the solvent. What is the molar mass of the unknown substance?


3.You add 1 kg of ethylene glycol (C2H6O2) antifreeze to 4500g of H2O in your car’s radiator. What would be the freezing point of the solution?


4.How many grams of Lactic acid (C6H10O5) should be dissolved in 650g of cyclohexane (C6H12) to raise the boiling point to 84.910C? The normal boiling point of cyclohexane is 810C and its molal boiling point is 2.75 0C/m.


5.The osmotic pressure of 250mL of an aqueous solution containing 0.150g of the enzyme lysozyme is found to be 0.0013atm at 24.850 Calculate the molar mass of lysozyme.


1
Expert's answer
2022-02-14T13:04:02-0500

a)

Molar mass of sucrose = 342g/mol


Moles of sucrose "= \frac{15}{342}=0.04386"


Molar mass of ethanol= 46g/mol

Moles of ethanol "=\frac{100}{46}=2.1739"


Total moles "= 2.1739+0.4386=2.21776"

Mole fraction of ethanol

"=\frac{2.1739}{2.21776}=0.98022"


"58=0.98022\u00d7P_{ethanol}"


"P_{ethanol}= 59.17mmHg"